OpenStudy (australopithecus):

Determine whether the sequence converges or diverges, then compute the limit {(2n-1)!/(2n+1)!}

OpenStudy (australopithecus):

I have never dealt with factorials before can anyone help me solve this?

OpenStudy (turingtest):

can we use the ratio test, or is that only for series?

OpenStudy (zarkon):

that is just for series

OpenStudy (turingtest):

thought so...

OpenStudy (zarkon):

just simplify and take the limit

OpenStudy (foolaroundmath):

\(n! = 1.2.3...n = (1.2.3.....(n-1)).n = n.(n-1)!\) This is a fundamental property of factorials. Now, the sequence that we have is \[a_{n} = \frac{(2n-1)!}{(2n+1)!}=\frac{(2n-1)!}{(2n-1)!.2n.(2n+1)} = \frac{1}{2n(2n+1)}\] can you take it off from here?

OpenStudy (turingtest):

ah, I see it now :0